A car and a bus started out from the same place, traveling in opposite directions. The average speed of the car was 15 miles per hour faster than the bus. After 5 hours, they were 535 miles apart. How fast was the car going?
Answer by Paul(988)

What you basically have to do is find the speed.
Since you are given distance, time and supportment of speed.
Equation will look like this Distance = time x speed.

Speed of the bus = x
Speed of the car = x+15
After 5 hours = 5 times the 2 components.
Total distance = 535

5(x+x+15)=535
10x+75=535
10x=460
x=46
x+15=46+15=61

Hence the speed of the bus was 46mph and the speed of the car was 61mph.